# Fast, systematic and robust relative binding free energies for simple and complex transformations : dual-LAO

### v1

@@ 2.1 Enhanced Alchemical Sampling: Dynamic $\lambda$ and Lambda-ABF-OPES
  • OPES-Explore (OPESe) : The Lambda-ABF scheme does not explicitly encompass capabilities to overcome energetic barriers orthogonal to the chosen collective variable ($\lambda$). This is addressed by the addition of the exploratory version of the On-the-fly Probability Enhanced Sampling (OPES) Invernizzi2020, Invernizzi2022 method, known as OPESe, a metadynamics-based method that aims at enhancing configurational sampling by lowering these orthogonal barriers.
@@ 2.4 Dual DBC Restraint Scheme
To counteract positional drift during the intermediate $\lambda$ states, we apply a flat-bottom harmonic restraint on the DBC of each ligand, using two independent, $\lambda$-dependent target distances, $\mathbf{d_{\mathrm{target1}}}$ and $\mathbf{d_{\mathrm{target2}}}$, as shown in Fig. 1(b).
@@
Restraints are maintained at $D_{\text{min}}$, but are gradually released near the end point corresponding to the decoupled ligand of interest by increasing the target distance to $D_{\text{max}}$. As explained in a further section, the value of $D_{\text{min}}$ is chosen so that the binding mode of the fully coupled ligand of interest is not biased. This ensures that the final (fully coupled and fully decoupled) states are unbiased.
Figure 1: Scaling of the intermolecular interactions and DBC restraint as a function of $\lambda$. (a) Scaling of the intermolecular interactions for ligand $L_{1}$ ($\lambda_{\text{ELE}}^{L_{1}}$, $\lambda_{\text{VDW}}^{L_{1}}$) and ligand $L_{2}$ ($\lambda_{\text{ELE}}^{L_{2}}$, $\lambda_{\text{VDW}}^{L_{2}}$) as a function of $\lambda$. (b) Dual-DBC target distance profiles ($\mathbf{d_{\mathrm{target1}}}$ and $\mathbf{d_{\mathrm{target2}}}$) applied as harmonic positional restraints on the distance between centers ($\mathrm{DBC}$) for complex alchemical transformations.
@@ 2.7 DBC Selection
To ensure that the DBC restraint does not introduce artificial bias during the alchemical simulation, we implemented a dual-DBC restraint strategy as explained earlier. Specifically, a minimum DBC value ($\text{DBC}_{\text{min}}$) was set at $2$ Å. This $\text{DBC}_{\text{min}}$ value was chosen to be lower than the upper bound of DBC coordinate value observed over the plain MD production run for each ligand, which was monitored using the Colvars library. A flat-bottomed harmonic restraint was applied to the DBC coordinate of both $L_{1}$ and $L_{2}$ during the dual-LAO simulations, following the DBC restraint schedule. This restraint employed a high force constant of $100$ $\text{kcal/mol/\AA }^{2}$ and was active only when the calculated DBC value exceeded the predetermined DBC cutoff.

### v2

@@ 2.1 Enhanced Alchemical Sampling: Dynamic $\lambda$ and Lambda-ABF-OPES
  • OPES-Explore (OPESe) : The Lambda-ABF scheme does not explicitly encompass capabilities to overcome energetic barriers orthogonal to the chosen collective variable ($\lambda$). This is addressed by the addition of the exploratory version of the On-the-fly Probability Enhanced Sampling (OPES) Invernizzi2020, Invernizzi2022 method, known as OPESe, a metadynamics-based method that aims at enhancing configurational sampling by adding a penalty to the already explored regions of the $\lambda$ interval.
@@ 2.4 Dual DBC Restraint Scheme
To counteract positional drift during the intermediate $\lambda$ states, we apply a flat-bottom harmonic restraint on the DBC of each ligand, using two independent, $\lambda$-dependent targets, $\mathbf{d_{\mathrm{target1}}}$ and $\mathbf{d_{\mathrm{target2}}}$, as shown in Fig. 1(b).
@@
Restraints are maintained at $D_{\text{min}}$, but are gradually released near the end point corresponding to the decoupled ligand of interest by increasing the target to $D_{\text{max}}$. As explained in a further section, the value of $D_{\text{min}}$ is chosen so that the binding mode of the fully coupled ligand of interest is not biased. This ensures that the final (fully coupled and fully decoupled) states are unbiased.
Figure 1: Scaling of the intermolecular interactions and DBC restraint as a function of $\lambda$. (a) Scaling of the intermolecular interactions for ligand $L_{1}$ ($\lambda_{\text{ELE}}^{L_{1}}$, $\lambda_{\text{VDW}}^{L_{1}}$) and ligand $L_{2}$ ($\lambda_{\text{ELE}}^{L_{2}}$, $\lambda_{\text{VDW}}^{L_{2}}$) as a function of $\lambda$. (b) Dual-DBC targets profiles ($\mathbf{d_{\mathrm{target1}}}$ and $\mathbf{d_{\mathrm{target2}}}$).
@@ 2.7 DBC Selection
To ensure that the DBC restraint does not introduce artificial bias during the alchemical simulation, we implemented a dual-DBC restraint strategy as explained earlier. Specifically, a minimum DBC value ($\text{DBC}_{\text{min}}$) was set at $2$ Å. This $\text{DBC}_{\text{min}}$ value was chosen to be higher than the upper bound of DBC coordinate value observed over the plain MD production run for each ligand, which was monitored using the Colvars library. A flat-bottomed harmonic restraint was applied to the DBC coordinate of both $L_{1}$ and $L_{2}$ during the dual-LAO simulations, following the DBC restraint schedule. This restraint employed a high force constant of $100$ $\text{kcal/mol/\AA }^{2}$ and was active only when the calculated DBC value exceeded the predetermined DBC cutoff.